OpenStudy (andu1854):

For what values of a and b is the following function is continuous at every x: -7 x ≤ -3 f(X) = ax-b -3< x < 1 3 x ≥ 1 Find the values is a and b ix f(x) continuous at every x?

OpenStudy (anonymous):

Let me try to help. According to the problem, we have: f(-3) = -7 f(1) = 3 Therefore,when we replace x with -3,then x with 1 into f(x) = ax-b, we have: -7 =-3a-b (#1) 3 =a-b (#2) we can solve this system of equations by the elimination method, for example. let's multiply both side of the second equation by -1(in order to eliminate the variable b) we have found a third equation: -3 = -a+b (#3) Now, by addition of the equations #1 and #3, we've got: -10 = -4a or 10 = 4a (by Multiplying both sides by -1) or a=10/4 or a=5/2 when we replace a = 5/2 into the equation (#3), we've found b=-1/2. The function can be written then: f(x) =( 5/2)x -(-1/2) or f(x) = (5/2)x + 1/2 I Hope I've help you.

OpenStudy (anonymous):

I hope I've helped you.

OpenStudy (andu1854):

Yes very much so,